OpenStudy (anonymous):

PLEASE HELP!!!!! WILL MEDAL AND FAN!!!!!!! Which is the extraneous solution of -x = √2x+15 (2x + 15 is under the √ sign. The entire right side of the equation is. Thanks)

OpenStudy (bloomlocke367):

what are your options?

OpenStudy (anonymous):

Oh, sorry. lol A. -5 B. -3 C. 3 D. 5

OpenStudy (bloomlocke367):

do you know what an extraneous solution is?

OpenStudy (anonymous):

Root of a changed equation I believe.

OpenStudy (bloomlocke367):

yes, and it DOES NOT solve for the original equation.

OpenStudy (anonymous):

I am pretty sure square both sides, then simplify to have x^2 - 2x - 15 = 0

OpenStudy (bloomlocke367):

yes, then find the solutions for x and whatever one does not work in the original is your answer

OpenStudy (anonymous):

then it would be something like (x + or - ?)(x + or - ?)

OpenStudy (bloomlocke367):

?.. what do you mean?

OpenStudy (anonymous):

?=some number, I am still working that one out. I learned that they need to be in two seperate parenthesis such as x ^2 + 8x + 15 would simplify to be (x + 5)(x + 3)

OpenStudy (bloomlocke367):

okay, to solve for x, using factoring like you're talking about, you need to find what 2 numbers multiply to get -15 but add up to get -2, since your equation is\[x^2-2x-15=0\]

OpenStudy (anonymous):

-5 and 3 (x - 5) (x + 3) Which then would say that the answer is either x = 5 or x = -3

OpenStudy (bloomlocke367):

yes! now plug those into the original, and see which one does not work.

OpenStudy (anonymous):

BECAUSE I subtracted/added the 5 and the 3 to 0 in seperate equations to get those two answers.

OpenStudy (anonymous):

-3 does not.

OpenStudy (bloomlocke367):

yep! good job :)

OpenStudy (anonymous):

Thanks! Here is the medal and fan....

OpenStudy (anonymous):

There you go. Thanks again!!

OpenStudy (bloomlocke367):

thank you cx

OpenStudy (bloomlocke367):

and no problem

OpenStudy (bloomlocke367):

wait... I think that's wrong...

OpenStudy (bloomlocke367):

\[-(-3)=\sqrt{2(-3)+15}\]\[3=\sqrt{-6+15}\]\[3=\sqrt 9\]

OpenStudy (anonymous):

Yes, it was wrong, but that is fine. :) I have a good enough grade in Algebra anyway. Thanks for the help though!